A bucket tied at the end of a 1.6 m long string is whirled in a vertical circle with a constant speed. What should be the minimum speed so that the water from the bucket does not spill when the bucket is at the highest position?

Grade:12

2 Answers

Srishti Sharma
6 years ago
For water not to spill out of bucket 
  mv^2/r = mg
  v = root rg = root 1.6 x root 10
  v = root 16 = 4 ms-1
Rituraj Singh
6 years ago
For a complete circular motion minimum u should be √5gl-mg3.2=1/2mv²-1/2mu²v²/2= 5gl/2 - 3.2gv²/2=40-32v²=16v=4m/s